OpenStudy (anonymous):

7(x -2) + 3(x + 2) = 5(x - 6)

OpenStudy (jadeishere):

7x - 14 + 3x + 6 = 5x - 20 10x - 8 = 5x - 30 5x = -22 x = -22/5 x = -4.4

OpenStudy (anonymous):

thank u

OpenStudy (jadeishere):

yeah, just add like terms and do PEMDAS

OpenStudy (anonymous):

ok n how i can get it in fraction

OpenStudy (jadeishere):

0.4 = 4/10 simplify that and you get 2/5 so -4 and 2/5

OpenStudy (anonymous):

ok gracias

OpenStudy (jadeishere):

De nada :)